Circular No. 9019

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA IN THE LARGE MAGELLANIC CLOUD 2009 W. Liller, Viña del Mar, Chile, reports the discovery of a possible

recurrent nova in the Large Magellanic Cloud on two unfiltered Kodak Technical Pan films taken on Feb. 5; at mag ≈ 10.6 on Feb. 5.067 UT, the star is located at α = 5^{h}40^{m}.7, δ = −66^{o}40′ (equinox 2000.0). Nothing brighter than mag 14.0 was visible at this position on Jan. 31.065, but it was confirmed to have faded to mag ≈ 11.8 on Feb. 7.092. This position is close to that given for the second 1971 LMC nova (α = 5^{h}40^{m}.6, δ = −66^{o}40′, equinox 2000.0; cf. IAUC 2353) and may be a recurrence of that nova; however, only one observation was reported for the earlier nova, and its V magnitude was given as 13.0, suggesting that it was well past maximum.

H. E. Bond, Space Telescope Science Institute; F. M. Walter, Stony Brook University; and E. Cosgrove and J. Espinoza, Cerro Tololo Interamerican Observatory (CTIO), report observations of Liller’s nova with the SMARTS 1.3- and 1.5-m telescopes at CTIO. A shallow I-band frame obtained with the 1.3-m telescope on Feb. 7.1 UT yields position α = 5^{h}40^{m}44^{s}.20, δ = −66^{o}40′11′′.6 (equinox 2000.0; ± 0′′.1 in each coordinate, measured relative to five nearby USNO-NOMAD-catalogue stars). A spectrum (range 385–455 nm; resolution 0.16 nm) obtained with the 1.5-m telescope on Feb. 8.1 shows broad, flat-topped emission at Hγ and Hδ with FWHM ∼ 3900 km/s, with P-Cyg wind absorption features at ∼ −3800 km/s. A similar, weaker feature due to He I 447.1-nm is also present.

COMET C/2007 N3 (LULIN) M. D. Hicks and K. J. Lawrence, Jet Propulsion Laboratory; and J.

M. Somers, Moorpark College, report that Bessel-R observations of C/2007 N3, obtained at the Palomar Hale 5-m telescope on Feb. 2.58 UT, yield magnitude R = 7.27 ± 0.02 in a 4′-diameter photometric aperture. This corresponds to an Afρ value (cf. IAUC 7342) of 16.6 ± 0.9 m, which is ≈ 0.1 that of post-perihelion measurements of comet 1P/Halley obtained at a similar heliocentric distance (Schleicher et al. 1998, Icarus 132, 397). Long-slit CCD spectrograms of C/2007 N3 were also obtained, and analysis is ongoing.

Visual total-magnitude and coma-diameter estimates: Jan. 4.25 UT, 7.1, 5′ (J. J. Gonzalez, Leon, Spain, 10×50 binoculars); 24.21, 6.9, 6′.5 (W. Hasubick, Buchloe, Germany, 10×50 binoculars); Feb. 4.50, 6.3, 10′ (C. S. Morris, Fillmore, CA, U.S.A., 10×50 binoculars); 7.86, 6.1, 11′ (K. Yoshimoto, Yamaguchi, Japan, 7×35 binoculars).

2009 February 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green